THE 5-MINUTE RESTAURANT GUIDE

One comes to Vancouver to eat - and to eat well. To make sure you have such an experience here’s a guide to restaurants within a 5-minute walk of the Harbour Centre campus we (Kathy Heinrich and Brian Alspach) think deserve mention. They are all marked on the attached map. An asterisk indicates that a copy of the menu is available in the folder at the registration desk. They are listed alphabetically.

1. *BISTRO! BISTRO! 16 Water Street, 682-2162. It was from here that we began our search of restaurants and also where we ended for dinner. Dinner was very good: a reasonable selection of wine by the glass and the oyster mushrooms with asiago cheese was delicious. Kitchen opens into the restaurant. Nice high-ceilings and a good atmosphere. Live jazz every Thursday and Sunday evening 6:30 - 10:30. Reservations recommended (in fact by 7pm on the Saturday evening we were there, people were being turned away). Note that prices on the menu we obtained will decrease in December!!

2. *CAFE PACIFICA Pan Pacific Hotel. Large open restaurant serving breakfast (including Japanese), lunch and dinner. Mostly Italian with Italian opera buffet ($34) on Friday evening. Lo-cal and vegetarian items noted on menu. Brunch accompanied by mellow jazz trio. Good views of the harbour.

3. CAMBIE HOTEL 312 Cambie, 683-7961 Lots of beer on tap - including Rainforest Amber (with such a name it can only be good). Pool tables, sporting events on TV, serves breakfast all day. Don’t think we’d eat here but an interesting place if you like the “wild side” of the city.

4. *CARRALL STREET CAFE 217 Carrall Street, 602-0225. Loved the art of the walls. Bar on the main floor with eating areas upstairs and at the back (where it’s glass roof, cobblestones and ivy - very pleasant). Occasional live light jazz - usually Wednesday but at other times as well. Menu looks interesting (e.g., duck breast with sour cherries and drambuie). 26 good wines available by the glass. Beer on tap. No dinner on Sunday.

5. CORDOVA CAFE 307 W Cordova, 683-5637. Lunch only. Blackboard menu. The most wonderful art (well, maybe you won’t agree) is on display on the walls and the floor. We didn’t eat here but saw a friend just about to order. Here’s her review: I recommend the food - tasty, vegetarian meals, generous portions. Lunch at around $5.00. Live, mellow music from the creative furniture corner.

6. *THE FIVE SAILS RESTAURANT Pan Pacific Hotel, 662-8112. Reportedly one of the best restaurants in the city although we haven’t eaten here yet. Without question one of the best water views - harbour, ships, North Vancouver. Plays good jazz and blues and has a bar. Expect dinner to take the evening and to pay more than at most places. A couple of samples from the menu: Smoked BC salmon with potato pancake and creme fraiche; caramelized swordfish steak. Didn’t check the wine menu.

7. *FORSTERS 200 Granville Street, 689-8888. Located inside the mall just across the road (towards the water) from the Waterfront. Advertises its specialty as Prime Rib but we noted the live Dungeness crab - best crab you’ll find anywhere. Very large restaurant overlooking the the Trade Centre, the harbour and the cruise ship terminal. Wine list moderate but a good selection. Deserts are kept in tall glass cases. Serves breakfast.

8. *IMPERIAL CHINESE SEAFOOD RESTAURANT 355 Burrard Street, 688-8191. Large menu but not your traditionally cheap Chinese food. Have eaten here and the food is excellent. (Could Mick Jagger be wrong!)

9. *INDIA VILLAGE RESTAURANT. 308 Water Street, 681-0678. Small and cosy with delicious spicy aromas. Small wine list of fine quality.

10. JOLLY TAXPAYER PUB AND BISTRO 828 West Hastings, 681-3574. Open weekdays from 11 am. Good standard pub food—wings, fettucine, fish and chips. We like the atmosphere.

11. *KISHU JAPANESE HOUSE 608 West Pender, 689-8883. In 1994 written up in the local paper as good value for money. Not a great decor (television on and lots of handwritten menus on the walls) but they say the food is good - at least that’s what the diners told us. Crowded for lunch during the week. Very cheap. We’d try it.

12. *KILIMANJARO 332 Water Street, 681-9913. One of our forever favorites. East African with lots of spice (and that just describes the owner). The appetizer plate is great. Specialty is ostrich, though we confess we’ve never ordered it. Also offers crocodile, wild boar and venison - no we haven’t eaten them either. Wonderful highceilinged room with atmosphere. Entertainment every Friday night from light jazz to African drumming. Small wine list but excellent quality. Below the restaurant is a coffee house - same menu but doesn’t have the atmosphere. For Sunday lunch, the upstairs is not open.

13. *LA LUNA CAFE 117 Water Street, 687-5862. Small. Art Deco — all wood, mirrors and metal. Make all their own food and serve low-fat. Offer breakfast (continental) and are open from 7am to 5pm every day. A daily special. We liked the people.

14. *PAPRIKA COTTAGE 106-12 Water Street, 684-7115. Small cafe serving lunch. Shaftsbury Cream Ale and Rainforest Amber on tap. Mixed oriental menu and sandwiches. Also serves cider and a house wine (and everyone should try BC cider!). Nice art on the walls.

15. *PALOMA 108 Cambie Street, 684-0688. This Polynesian bar and grill also serves fettucine (because they’re in Gastown). Menu in three languages. Interesting tropical cocktails.

16. *RAINTREE RESTAURANT AT THE LANDING 375 Water Street, 588-5570. We ate here several years ago before it moved to this location. Loved it then and reports are it’s as good as ever. (For those from Victoria the owner also owned Harvest Moon Cafe — you’ll recognise the flying cows.) Many consider it one of Vancouver’s best, View over the Harbour (well, over the harbour over the railway tracks). An interesting West Coast (Washington, Oregon, California, BC) wine list. Serves 2 or 3 micro-brews. Great atmosphere - and great wall hangings by Rozz Marshall (buy them at the Bushlen Mowatt Gallery on Georgia).

17. RIO RIO 102 Water, 685-1145. Tapas bar. Order lots and lots of small dishes and share them with, of course, sangria. Had dinner here about a year ago and thought the food was pretty good.

18. *ROGUES CAFE 302 W. Cordova, 683-0808. Got a lot of points during our visit by playing Billie Holiday as background music. Excellent food (or so we were told by the owner). Very small wine list but it was a good selection.

19. SITAR 8 Powell Street, 687-0049. We like the atmosphere (both the feeling of the rooms and the smells of spices). Ate here a lot many years ago. Lunch and dinner served daily.

20. SODAS DINER 375 Water Street, 222-9922. Pretty good burger served as in a pretty good diner. Microbrews on draft (Granville Island and Shaftesbury). Serves BC Cider. Specialty drinks are “adult” shakes!

21. STARBUCKS several locations. Considered by many the place to drink coffee (and they have good hot chocolate). It’s hard to travel more than a few blocks anywhere in Vancouver without finding yourself at Starbucks!!

22. *STEAMWORKS BREWING CO. 375 Water Street, 689-2739. And yes, they do - specializing in ales. Sometimes carry other microbreweries. Another specialty is single malt scotch - and there’s an interesting wine collection. Before we forget, food is also served - although the menu we obtained has since changed - so expect new items. Long bar downstairs (held up by gargoyles) and two pool tables. There’s an exciting noisy feel about the place.

23. *UMBERTO AL PORTO 321 Water Street, 683-8376. One of several Umberto restaurants in Vancouver - all with excellent reputations. We have eaten here and it is no exception. The entrance is down the alley off Water Street (the menu is posted on the corner). Nice decor and a view of the harbour (yes, over the railway tracks). No lunch on Saturday and closed on Sunday.

24. *WATER ST. CAFE 300 Water Street, 689-2832. Two walls of glass let you sit and watch the tourists go by (and there are lots of them!). Pleasant atmosphere with some nice art on the walls. We were told the food was excellent. Wine list moderate with prices from $20 to $40.